Business Process Management Systems (2II55) ”Drankorgel”

Oliver Schinagl (0580852) Maik Teurlings (0591600) Wilrik De Loose (0601583)

March 31, 2008

1 Settings

1.1 Organizational context

The Party store is composed of four groups - the warehouse, store, administration and logistics department. The organizational model of the party store is presented below where blue color represents groups and the green color represents roles.

Name Executor Description

Order

Store employee The store employee takes an order of a customer.

Receive cash Store employee The store employee receives the cash money from the

customer.

Create bill Administrative employee The order is processed and the bill for the customer is being created.

Send bill Administrative employee

The bill is being send to the customer.

Payment re- ceived Administrative manager

The manager marks the order as being payed.

Payment timeout System The manager sends a new bill if the bill has not been payed for three weeks.

Plan logis- tics Logistics employee The employee plans the logistics used for the shipment of the ordered goods.

Pick order Warehouse employee

The employee picks the needed goods for the order.

Order goods Warehouse manager The manager orders the goods needed to complete the order.

Receive goods Warehouse employee The employee e_{1} receives the goods ordered by the manager to complete the order.

Inspect order Warehouse employee The employee e_{2}, where e_{1} ≠ e_{2}, inspects the order for completeness.

Cancel order Manager The manager can decide to cancel an order